# Project Page: https://gyhdog99.github.io/projects/mocle/

### v4

@@ 4.4 Ablation Studies
In the proposed MoCLE method, the value of temperature plays an important role in controlling the contribution of the universal expert. Specifically, as shown in Eq. (1), $\tau$ controls the sharpness of the gate distribution, while the output of the universal expert is weighted by $1-G_{\text{max}}$. Therefore, as $\tau$ decreases, $G_{\text{max}}$ increases, and finally the contribution of the universal expert decreases. As shown in Table 7, the results are consistent with our understandings. When a $\tau$ is too small, i.e., 0.01, the model generalizes worse to unseen tasks like SciQA and VSR. Similarly, when $\tau$ is too large, i.e., $0.2$, the role of task experts is diminished, leading to task conflicts. The temperatures of $0.05$ and $0.1$ seem to achieve a balance between specialization and generalization of the model, and we use 0.05 as the default setting for the temperature.
@@ 4.5 Visualizations
Figure 3: (a) Clustering assignment of the training datasets when $K=64$. The labels on the vertical axis indicate the names of the datasets. Within each labeled block, multiple rows represent different subsets, each of which is associated with a unique task template. The horizontal axis denotes the cluster index to which these subsets are assigned. (b) t-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
4.5.1 Clustering.
@@
We first show the justification to represent the training data via their instructions. Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model [49]. After that, we visualize the data in Figure 2(b) via t-SNE [44]. According to Figure 2(b), we observe that (1) samples from different datasets but from the same task are grouped together. For example, coco_vqg, aok_vqg, and ok_vqg are all VQG (visual question generation) tasks (triangle markers) and reside on the left part of the figure. (2) samples from similar tasks are closed to each other. For example, coco_cap and textcaps both belong to the image captioning (CAP) task (small dots) and stay close to each other. Similarly, both the visual question answering (VQA, cross markers) task with datasets (e.g., aok_vqa, coco_vqa, ocrvqa, and ok_vqa) and conversation (CONV) task with datasets (e.g., llava_conversation, llava_detail, and llava_reason) involve answering user questions, which lie in the middle part of Figure 2(b), suggesting that instructions are good representatives of the training data.
@@
We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering. Figure 2(a) shows the cluster assignment of the training data. Here, each row in the heatmap denotes a subset of a dataset. The subset is obtained by applying the task template (Sec. 3.1) on the samples of the dataset. We observe (1) different subsets of the same datasets are assigned to similar clusters. For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters, while aok_vqg, coco_vqg, and ok_vqg mainly occupy 4 other clusters. (2) datasets of similar tasks are assigned to common clusters. For example, llava_150k including llava_detail, llava_detail and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions. Those evidences justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
4.5.2 Routing results.
@@
Figure 4 visualizes the routing decisions of our MoCLE and Sentence-MoLE. We obtain both results from one mixture of LoRA, i.e., one linear module in a layer. The routing results are aggregated by the subset of datasets similar to Figure 2(a). Shown in Figure 3(a), MoCLE could achieve task-level routing for the inputs. For example, datasets from VQA and VQG tasks are handled by expert 0 and 3 respectively. In addition, VQA expert (i.e., expert 0, which is mostly occupied by VQA task) is fundamental for most of the datasets. That is, most held-out datasets activate the VQA expert. This could be justified by the fact that most vision-language tasks take the form of question answering. However, this does not indicate that the entire model collapses to a single expert because knowledge from other tasks are learned by the universal expert and are leveraged during evaluation. Further, the remaining 3 experts are not dispensable as without them all tasks would clutter in a single expert, resulting in task conflicts. Instead, routing pattern of Sentence-MoLE in Figure 3(b) reveals little correlations between tasks/datasets and experts. That is, different datasets/tasks obtain similar routing decisions, still suffering from task conflicts.
@@
Figure 4: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE. The setup of the vertical axis is similar to Figure 2(a) except that we also include the held-out tasks. They are separated by a dotted line on the vertical axis. The horizontal axis corresponds to the index of the LoRA experts.
@@ 5 Conclusions
In this paper, we first show that the task conflicts exists in vision language instruction tuning through extensive experiments. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on 11 unseen tasks show the effectiveness of our MoCLE.
Acknowledgement
@@
Appendix
Appendix 0.A Sources of Results in Table 1
For LLaVA (13B) and MiniGPT-4 (13B), the results on GQA, VSR, OKVQA and IconQA are from the Table 3 of [5], the results on POPE and A-OKVQA (Direct) are from Table 3 and 6 of [33], respectively. In particular, the performance of MiniGPT-4 (13B) on MME is from its official leaderboard  [^1] https://github.com/BradyFU/Awesome-Multimodal-Large-Language-Models/tree/Evaluation. All the task performances of InstructBLIP (7B) are obtained from its reproduced version implemented in this paper.
Appendix 0.B Training Details
Following [14], we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec. 1. We train those models with a maximum of 60K steps and a batch size of 128. The AdamW optimizer [26] is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05. We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
Appendix 0.C Weights of the Universal Experts
@@
Appendix 0.D Datasets for Vision-language Instruction Tuning
@@
Here we give a detailed description of datasets used during instruction tuning. As shown in Table 9, we use 13 datasets in total. Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix 0.E). For LLaVA-Instruct150K [38], we do not apply any task template as it has been well formatted. For Web Capfilt, we use COCO+VG+CC+SBU (14M images) as [29]. Regarding with which split is used, please refer to the InstructBLIP paper [14] for more details. During training, we sample datasets using the data balancing strategy as in [14].
Appendix 0.E Task Templates
@@
Following [14], we use the same set of task templates for instruction tuning and held-in/out evaluation. Please refer to Tables 14 and 15 for training and evaluation templates.
Appendix 0.F Case Studies
0.F.1 Conversations
In Table 10, we instruct the model to conduct a very difficult object counting task. The correct answer for this question is 63, which is quite hard for existing LVLMs. InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a much more proper manner.
Similarly, in Table 13, we ask the model to describe the image in a detailed manner. InstructBLIP still mistakes this query as an image caption task and gives very short caption to this image. Instead, with MoCLE, the model correctly understands the “in details” in the instruction and provides sufficient details.
Datasets Data Source Task Template Group
aok_vqa A-OKVQA [52] VQAMC
aok_vqg A-OKVQA [52] VQG
coco_cap COCO [35] CAP
coco_vqa VQAv2 [19] VQA
coco_vqg VQAv2 [19] VQG
ocrvqa OCR-VQA [46] VQA
ok_vqa OKVQA [45] VQA
ok_vqg OKVQA [45] VQA
textcaps TextCaps [55] OCRCAPS
capfilt Web CapFilt [29] CAP
llava_conversation LLaVA-Instruct150K [38] -
llava_detail LLaVA-Instruct150K [38] -
llava_reason LLaVA-Instruct150K [38] -
Table 9: List of datasets used for training.
Model responses to counting-related queries. [Uncaptioned image] User How many books are there? New Bing I see that you have sent me an image. According to the description you provided, there are 30 books on the bookshelf. InstructBLIP 100 +MoCLE There are a lot of books on the shelves. It is difficult to say exactly how many books there are, but it appears to be a large number. Table 10: Model responses to counting-related queries.
Model responses to OCR-related queries. [Uncaptioned image] User What are all the exact scene texts in the image? New Bing According to the description you provided, the image contains a license plate with the text “S3302CD”. If you have any other questions or if there is anything else I can help you with, please let me know! InstructBLIP The exact characters in the image are “S 3000 CD.” +MoCLE The exact characters in the image are “S 3302 CD.” Table 11: Model responses to OCR-related queries.
Model responses when asked to introduce a famous person. [Uncaptioned image] User Can you give an introduction to this person? New Bing I’m sorry, I cannot give an introduction to this person as I am not allowed to describe people’s identity, gender, race or emotions InstructBLIP albert einstein +MoCLE Albert Einstein was a German-born theoretical physicist who developed the theory of relativity. He is considered one of the most influential scientists of the 20th century. Table 12: Model responses to introduce a celebrity.
0.F.2 Sample Instructions in Clusters
In Table 16, we showcase some sample instructions assigned to different clusters. Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting, justifying the usage of a large number of instruction clusters. Moreover, in Table 7, we show that the model performs poorly with only a few clusters, i.e., $K=4$, and gets better as we increase $K$. With a few clusters, VQA, VQG, CONV and CAP tasks will occupy 4 clusters, leaving no more space for samples with different topics. Thus, the model might perform worse when mixing these topics.
Model responses when queried to give detailed image descriptions. [Uncaptioned image] User Introduce me this image in details. New Bing It appears to be a photograph of an adult elephant walking on a beach. The elephant is walking towards the camera and is in the center of the frame. The background consists of the ocean and the sky. The elephant is walking on the sand and there are a few birds in the foreground. The elephant appears to be dark grey in color and has large ears and tusks InstructBLIP an elephant walking along the beach +MoCLE An elephant is walking along the beach, with birds in the background. The elephant’s trunk is extended towards the water, and it appears to be enjoying its time at the beach. Table 13: Model responses to give detailed image description.
Template Group
Task Template
CAP
“A short image caption:”
“A short image description:”
“A photo of”
“An image that shows”
“Write a short description for the image.”
“Write a description for the photo.”
“Provide a description of what is presented in the photo.”
“Briefly describe the content of the image.”
“Can you briefly explain what you see in the image?”
“Could you use a few words to describe what you perceive in the photo?”
“Please provide a short depiction of the picture.”
“Using language, provide a short account of the image.”
“Use a few words to illustrate what is happening in the picture.”
VQA
“{Question}”
“Question: {Question}”
“Question A short answer to the question is”
“Q: {Question} A:”
“Question: {Question} Short answer:”
“Given the image, answer the following question with no more than three words. {Question}”
“Based on the image, respond to this question with a short answer: {Question}. Answer:”
“Use the provided image to answer the question: {Question} Provide your answer as short as possible:”
“What is the answer to the following question? {Question}”
“The question {Question} can be answered using the image. A short answer is”
VQG
“Given the image, generate a question whose answer is: {Answer}. Question:”
“Based on the image, provide a question with the answer: {Answer}. Question:”
“Given the visual representation, create a question for which the answer is {Answer}.”
“From the image provided, craft a question that leads to the reply: {Answer}. Question:”
“Considering the picture, come up with a question where the answer is: {Answer}.”
“Taking the image into account, generate an question that has the answer: {Answer}. Question:”
@@ 0.F.2 Sample Instructions in Clusters
Evaluation Dataset
Instruction
GQA, iVQA,
“Question: {} Short answer:”
@@
A-OKVQA(MC),
“Question: {} Options: {}. Short answer:”
@@
NoCaps, Flickr30k
“A short image description:”
TextVQA
“OCR tokens: {}. Question: {} Short answer:”
ScienceQA
“Context: {} Question: {} Options: {}. Answer:”
HatefulMemes
“This is an image with: “{}” written on it. Is it hateful? Answer:”
VSR
“Based on the image, is this statement true or false? “{}” Answer:”
Visdual Dialog
@@
Cluster
Instruction Samples
Topics
1
“Q: what is being done to the food in the glass fronted box? A:”
Food

### v5

@@ 4.4.5 Temperature
In the proposed MoCLE, the temperature plays an important role in controlling the contribution of the universal expert. Specifically, as shown in Eq. (1), $\tau$ controls the sharpness of the gate distribution, while the output of the universal expert is weighted by $1-G_{\text{max}}$. Therefore, as $\tau$ decreases, $G_{\text{max}}$ increases, and finally the contribution of the universal expert decreases. As shown in Figure 3, the results are consistent with our understandings. When $\tau$ is either too small (0.01) or large (0.2) can lead to inferior results. The temperatures of $0.05$ and $0.1$ seem to achieve a balance between specialization and generalization of the model. In the experiments, we use temperature of 0.05 as default.
4.4.6 Number of Task Experts
As demonstrated in Figure 3, more task experts usually provides with stronger capacity. Specifically, when only 2 task experts are employed, we observe inferior overall results. This model has similar capacity to the single LoRA model in Sec. 4.4.1, where only one LoRA encounters difficulties in fitting a diverse set of tasks. When the number of task experts is increased to 4, the performance gets improved. When the number of task experts becomes 8, it behaves similarly to the 4-expert case, which indicates that the benefit of increasing capacity converges as we use more task experts. Hence, we use 4 task experts as the default setting.
@@ 4.5 Visualizations
Figure 4: T-SNE visualization of the instruction encoding. Different colors correspond to different datasets, while the shape of the markers indicates the task category defined manually.
Figure 5: Clustering assignment of the training datasets when $K=64$. The labels on the y-axis indicate the names of the datasets. The x-axis denotes the cluster index to which the subsets are assigned.
4.5.1 Clustering
@@
We first show the justification to represent the training data via their instructions. Specifically, for each dataset, we sample 100 examples and encode their instructions with the all-MiniLM-L6-v2 variant of the Sentence Transformer model Reimers and Gurevych (2019). We then visualize the data in Figure 4 via t-SNE Van der Maaten and Hinton (2008). As can be seen, (1) Samples from the same task are grouped together. For example, all visual question generation (VQG, triangle markers) data reside on the left part of the figure. (2) Samples from similar tasks are close to each other, e.g., coco_cap and textcaps both belong to the image captioning (CAP, small dots) task and stay close to each other at the lower right of the figure. Similarly, both visual question answering (VQA, cross markers) and conversation (CONV, y-shape) data involve answering user questions, which lie in the middle part of the figure, suggesting that instructions are good representatives of training data.
@@
We then cluster all the instructions of the examples in the training data into 64 groups using $k$-means clustering. Figure 5 shows the cluster assignment of the training data. Here, each row in the heatmap denotes a subset of a dataset. The subset is obtained by applying the task template (Sec. 3.1) on the samples of the dataset. We observe the following: (1) Different subsets of the same datasets are assigned to similar clusters. For example, aok_vqa, coco_vqa, and ok_vqa are in the first several clusters. (2) Datasets of similar tasks are assigned to common clusters. For example, llava_150k including llava_detail, llava_reason and llava_conversation and a series of VQA tasks share the first several clusters as they are to answer questions. These justify the use of clustering on task instructions as an automatic partition strategy for training datasets.
4.5.2 Routing Results
@@
Figure 6 visualizes the routing decisions of the proposed MoCLE and Sentence-MoLE. We obtain both results from one mixture of LoRA, i.e., one linear module in a layer. The routing results are aggregated by the subset of datasets similar to Figure 5. As can be seen from Figure 6(a), MoCLE can achieve task-level routing for the inputs. For example, datasets from VQA and VQG tasks are handled by expert 0 and 3, respectively. Instead, routing pattern of Sentence-MoLE in Figure 6(b) reveals little correlations between datasets and experts. That is, different datasets obtain similar routing decisions, and thus still suffer from task conflicts.
@@
Figure 6: Routing decisions of one LoRA mixture for MoCLE and Sentence-MoLE. The setup of the vertical axis is similar to Figure 5 except that we also include the held-out tasks. They are separated by a dotted line on the vertical axis. The horizontal axis corresponds to the index of the LoRA experts.
@@ 5 Conclusions
In this paper, we first show through extensive experiments that task conflicts exist in vision language instruction tuning. To address this, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel MoE architecture designed to activate the task-customized model parameters based on the instruction clusters. In addition, we achieve task specialization and generalization in MoCLE simultaneously via a separate universal expert. Comprehensive evaluations of MoCLE on both held-out/in tasks show the effectiveness of MoCLE.
6 Limitations
Although effective, we mainly focus on task conflicts among text-based conversation tasks in this paper, while the support of our MoCLE for more complicated visual perception tasks is appealing, which has shown more severe task conflicts with the conversation tasks Zhu et al. (2022).
7 Acknowledgement
@@
Appendix A Training Details
A.1 InstructBLIP
Following Dai et al. (2023), we adopt the same training configurations for the mentioned models such as the proposed MoCLE, the reproduced InstructBLIP (7B) and the task experts in Sec. 1. We train those models with a maximum of 60K steps and a batch size of 128. The AdamW optimizer Kingma and Ba (2014) is used, with $\beta_{1}$ as 0.9, $\beta_{2}$ as 0.999, and a weight decay as 0.05. We apply a linear warmup of the learning rate during the initial 1000 steps, increasing from ${10}^{-8}$ to ${10}^{-5}$, followed by a cosine decay with a minimum learning rate of 0.
A.2 LLaVA-1.5
We follow Liu et al. (2023a) for the training configuration. Specifically, for LLaVA-150KLiu et al. (2023a), Geo170KGao et al. (2023a) and Med. Mix, i.e., VQA-RADLau et al. (2018), SLAKELiu et al. (2021) and Path-VQAHe et al. (2020), we train the model for 1, 2 and 9 epochs, respectively. When training on all of these datasets, we copy each dataset $k$ times ($k$ is the number of epochs it is trained independently) and merge them into a single dataset. For each training job, we use a batch size of 128, weight decay of 0 and learning rate of $41e-$ 4, which is warmed up from 0 during the initial 3% steps and followed by a cosine decay with a minimum learning rate of 0.
Appendix B Weights of the Universal Experts
@@
Appendix C Data
We list the training and evaluation data for InstructBLIP and LLaVA-1.5 in Table 8.
C.1 Dataset Abbreviation
@@
For InstructBLIP, we further show in Table 9 (1) the abbreviation used in Sec. 4.5 for each dataset and (2) their manually defined task category. As shown in Table 9, we use 13 datasets in total. Here multiple datasets might be associated with the same data sources because these sources are formatted by different groups of task templates (see Appendix D). For LLaVA-150K Liu et al. (2023b), we do not apply any task template as it has been well formatted.
Models Training datasets Evaluation Datasets InstructBLIP Flickr30K Young et al. (2014), GQA Hudson and Manning (2019) Web CapFilt Li et al. (2023a) VSR Liu et al. (2022a), IconQA Lu et al. (2021) A-OKVQA Schwenk et al. (2022) TextVQA Singh et al. (2019), Hateful Memes Kiela et al. (2020) TextCaps Sidorov et al. (2020), VQAv2Goyal et al. (2017) ScienceQA Lu et al. (2022), MSVD-QA Xu et al. (2017) OKVQA Marino et al. (2019), COCO Lin et al. (2014) MSRVTT-QA Xu et al. (2017), iVQA Yang et al. (2021) OCRVQA Mishra et al. (2019), LLaVA-150K Liu et al. (2023b) MME Fu et al. (2023), POPE Li et al. (2023d) OKVQA, A-OKVQA, VQAv2 LLaVA-1.5 LLaVA-665K Liu et al. (2023a) MME, MMBenchLiu et al. (2023c), ScienceQA Geo170K Gao et al. (2023a), VQA-RAD Lau et al. (2018) GeoQA Chen et al. (2021a), VQA-RAD SLAKE Liu et al. (2021), PathVQA He et al. (2020) SLAKE, PathVQA
Table 8: Datasets used for training and evaluation. *: the train split of this dataset is used during instruction tuning.
Datasets Data Source Task Template Group aok_vqa A-OKVQA VQAMC aok_vqg A-OKVQA VQG coco_cap COCO CAP coco_vqa VQAv2 VQA coco_vqg VQAv2 VQG ocrvqa OCR-VQA VQA ok_vqa OKVQA VQA ok_vqg OKVQA VQA textcaps TextCaps OCRCAPS capfilt Web CapFilt CAP llava_conversation LLaVA-150K - llava_detail LLaVA-150K - llava_reason LLaVA-150K -
Table 9: Abbreviation and manually defined task categories for the training datasets of InstructBLIP.
Appendix D Task Templates
@@
For InstructBLIP, we use the same set of task templates following Dai et al. (2023) for instruction tuning and held-in/out evaluation. Please refer to Tables 10 and 11 for training and evaluation templates
Template Group Task Template CAP A short image caption: A short image description: A photo of An image that shows Write a short description for the image. Write a description for the photo. Provide a description of what is presented in the photo. Briefly describe the content of the image. Can you briefly explain what you see in the image? Could you use a few words to describe what you perceive in the photo? Please provide a short depiction of the picture. Using language, provide a short account of the image. Use a few words to illustrate what is happening in the picture. VQA {Question} Question: {Question} Question A short answer to the question is Q: {Question} A: Question: {Question} Short answer: Given the image, answer the following question with no more than three words. {Question} Based on the image, respond to this question with a short answer: {Question}. Answer: Use the provided image to answer the question: {Question} Provide your answer as short as possible: What is the answer to the following question? “{Question}” The question “{Question}” can be answered using the image. A short answer is VQG Given the image, generate a question whose answer is: {Answer}. Question: Based on the image, provide a question with the answer: {Answer}. Question: Given the visual representation, create a question for which the answer is “{Answer}”. From the image provided, craft a question that leads to the reply: {Answer}. Question: Considering the picture, come up with a question where the answer is: {Answer}. Taking the image into account, generate an question that has the answer: {Answer}. Question:
@@
Evaluation Dataset Instruction
GQA, iVQA, Question: {} Short answer:
@@
A-OKVQA(MC), Question: {} Options: {}. Short answer:
@@
NoCaps, Flickr30k A short image description:
TextVQA OCR tokens: {}. Question: {} Short answer:
ScienceQA Context: {} Question: {} Options: {}. Answer:
HatefulMemes This is an image with: “{}” written on it. Is it hateful? Answer:
VSR Based on the image, is this statement true or false? “{}” Answer:
@@
Appendix E Case Studies
E.1 Conversations
In Table 12, we instruct the model to conduct a very difficult object counting task. The correct answer for this question is 63, which is quite hard for existing LVLMs. InstructBLIP fails to give the correct answers, while with MoCLE, InstructBLIP can respond the user query in a more proper manner.
Similarly, in Table 15, we ask the model to describe the image in a detailed manner. InstructBLIP still mistakes this query as an image caption task and gives a very short caption. Instead, with MoCLE, the model correctly understands the “in details” in the instruction and gives sufficient details.
Model responses to counting-related queries.
User How many books are there?
New Bing I see that you have sent me an image. According to the description you provided, there are 30 books on the bookshelf.
InstructBLIP 100
+MoCLE There are a lot of books on the shelves. It is difficult to say exactly how many books there are, but it appears to be a large number.
Table 12: Model responses to counting-related queries.
Model responses to OCR-related queries.
User What are all the exact scene texts in the image?
New Bing According to the description you provided, the image contains a license plate with the text “S3302CD”. If you have any other questions or if there is anything else I can help you with, please let me know!
InstructBLIP The exact characters in the image are “S 3000 CD.”
+MoCLE The exact characters in the image are “S 3302 CD.”
Table 13: Model responses to OCR-related queries.
Model responses when asked to introduce a famous person.
User Can you give an introduction to this person?
New Bing I’m sorry, I cannot give an introduction to this person as I am not allowed to describe people’s identity, gender, race or emotions
InstructBLIP albert einstein
+MoCLE Albert Einstein was a German-born theoretical physicist who developed the theory of relativity. He is considered one of the most influential scientists of the 20th century.
Table 14: Model responses to introduce a celebrity.
E.2 Sample Instructions in Clusters
In Table 16, we showcase some sample instructions assigned to different clusters. Though all the instructions in the 4 selected clusters belong to VQA-related tasks, they focus on various perspectives such as food, pet, men, and counting, justifying the usage of a large number of instruction clusters.
Appendix F Future Works
Although effective, we mainly focus on task conflicts among text-based conversation tasks (e.g., image captioning and VQA Gou et al. (2023)) in this paper, while the support of our MoCLE for more complicated visual perception tasks (e.g., object detection Han et al. (2021); Li et al. (2022)) is appealing, which has shown more severe task conflicts with conversations Zhu et al. (2022). Moreover, the applicability of MoCLE to visual pre-training Chen et al. (2021b, 2023b) and diffusion models Chen et al. (2023d); Gao et al. (2023b); Liu et al. (2023d); Li et al. (2023c) is also a potential research direction, considering the ubiquity of negative transfer phenomenon.
Model responses when queried to give detailed image descriptions.
User Introduce me this image in details.
New Bing It appears to be a photograph of an adult elephant walking on a beach. The elephant is walking towards the camera and is in the center of the frame. The background consists of the ocean and the sky. The elephant is walking on the sand and there are a few birds in the foreground. The elephant appears to be dark grey in color and has large ears and tusks
InstructBLIP an elephant walking along the beach
+MoCLE An elephant is walking along the beach, with birds in the background. The elephant’s trunk is extended towards the water, and it appears to be enjoying its time at the beach.
Table 15: Model responses to give detailed image description.
Cluster Instruction Samples Topics
1 “Q: what is being done to the food in the glass fronted box? A:” Food